OpenStudy (anonymous):

Initial Value Problem: dy/dx=y/ln y; y(0)=e I got it all wrong and used integration by parts. Please help...

OpenStudy (anonymous):

is it \[y'=\frac{y}{\ln y}\]

OpenStudy (anonymous):

y'=y/ln(y) dy/dx=y/ln(Y) ln(y)/ydy=dx

OpenStudy (anonymous):

NOw you can see the d/dy(ln(y) is sitting in there so its simply 1/2(ln(y))^2=x+C after integrating both sides (ln(y))^2=2x+C ln(y)=SQRT(2x+C) y=e^SQRT(2x+C) e=e^(sqrt(0+c) c=1 y=e^SQRT(2x+1) do you agree? james wanna help out?

OpenStudy (anonymous):

It seems steps were left out. From the beginning, what happened to the dx?

OpenStudy (anonymous):

ok want me to go line by line with you?

OpenStudy (anonymous):

Please, so frustrated here.

OpenStudy (anonymous):

haha alright jus relax these are confusing at 1st. ok y'=dy/dx do you agree

OpenStudy (anonymous):

If there was no dx, yes. But there is, so what happened to it?

OpenStudy (anonymous):

an example y=x^2 the y' or the derivative of y with respect to x =2x OR dy/dx=y's

OpenStudy (anonymous):

ok I guess it would make sense if it was d/dy=, rather than dy/dx. I don't see x's role.

OpenStudy (anonymous):

when you see y' program it in your brain to see dy/dx

OpenStudy (anonymous):

so dy/dx=y/ln(y) do you agree up to there

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok dy/dx=y/ln(y) dy(ln(y))=ydx yes? ln(y)/ydy=dx yes?

OpenStudy (anonymous):

now thiis should look familiar as a separable equation

OpenStudy (anonymous):

The x disappeared, and is back again?

OpenStudy (anonymous):

no, i simply multiplied the dx to the other side and divided by ln(y)/y to get all the dy and y's on one side and isolate the x

OpenStudy (anonymous):

isolated the DX

OpenStudy (anonymous):

ok I did that. Still confused how y' would equal dy/dx. Seems like it would equal just dy

OpenStudy (anonymous):

youll see in the end why its dy/dx and not just dy

OpenStudy (anonymous):

ok so we agree up to ln(y)/ydy=dx?

OpenStudy (anonymous):

So if you mult. bo sides by dx, wouldn't that leave you with just dy on left side?

OpenStudy (anonymous):

yes dy on the other side but then i multiplies the right side by the inverse of y/ln(y) so it would so away on the right and appear on the left with the dy as ln(Y)/y \[\int\limits_{}^{}\ln(y)/ydy=\int\limits_{}^{}dx\] then we integrate both sides

OpenStudy (anonymous):

ok. Did that, not sure how the x is magic, but ok.

OpenStudy (anonymous):

just take it as truth right now which it is. so can you tell me what the integral of both sides is?

OpenStudy (anonymous):

[ln y]^2/2=x+c

OpenStudy (anonymous):

CORRECT! now can you isolate the y?

OpenStudy (anonymous):

I just plugged 0 in for y for IVP and got 1/2=x+c

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

no you need to isolate y before you plug in your y(0)

OpenStudy (anonymous):

so 1/2(ln(y))^2=x+c (ln(y))^2=2x+C can you finish it?

OpenStudy (anonymous):

sqrt both sides, now I have sqrt(2x+c). Where does the IVP come into play here?

OpenStudy (anonymous):

Should I raise both sides to base e? Or am I ready to plug in 0 for y?

OpenStudy (anonymous):

no because you still have ln(y)=y so you need to e^SQRT(2x+C)

OpenStudy (anonymous):

ln(Y)=SQRT(2x+C) i meant

OpenStudy (anonymous):

NOW YOU HAVE y=e^SQRT(2x+C) now in calc one and 2 the teach said take the derivative of this function so you took the derivative of y w.r.t x dy/dx?

OpenStudy (anonymous):

I have y on left by iteself now. I plug in 0 for y, and e on right for IVP. I get 0=e to some power..impossible.

OpenStudy (anonymous):

yes so Y(x)=e^SQRT(2x+C) y(0)=e=e^sqrt(0+C) now solve for c

OpenStudy (anonymous):

y is on left side, so I put 0 in for it. I have e^something equalling the zero on left?

OpenStudy (anonymous):

no y(0)=e you have to have the e^SQRT(2x+C)=y(0) which is e

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

e=e^sqrt(0+C) take ln of both sides 1=SQRT(0+C) 1^2=C C=1 so now your final answer is y=e^SQRT(2x+1)

OpenStudy (anonymous):

How did you get 1 on the left side? It was y(0) before.

OpenStudy (anonymous):

I had e^ (ln y), so it turned into just y. Now it is just 1?

OpenStudy (anonymous):

your IVP states y(0)=e which means y(0) must =e on the right hand side so you need to solve for C to satisfy this

OpenStudy (anonymous):

so any IVP set you y(0)= to your solved eq and then solve for C

OpenStudy (anonymous):

because when you have y(x)=e^SQRT(2x+c) when x=0 you get e thats the point of it

OpenStudy (anonymous):

when c=1

OpenStudy (anonymous):

It looks good to me, thank you!!

OpenStudy (anonymous):

i am happy i helped you, have a great day

OpenStudy (anonymous):

You too. It's nice to see there are patient, friendly people out there